DCPAM Reference Manual

GFD Dennou Club

平成 17 年 2 月 18 日

目次

1 Module dycore hs94forcing mod 2 1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Known Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 Note . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.5 Future Plans . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.6 Dependency . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.7 Public Interface . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.8 Procedure Interface . . . . . . . . . . . . . . . . . . . . . . . 2

1.8.1 Initialize module and . . . . . . . . . . . . . . . . . . 2 1.8.2 Calculate Heating and Dissipation for ”Held and

Suarez(1994)” . . . . . . . . . . . . . . . . . . . . . . . 3 1.8.3 Terminate module . . . . . . . . . . . . . . . . . . . . 4

1 Module dycore hs94forcing mod

• Developers: Morikawa Yasuhiro

• Version: $Id: dycore hs94forcing.f90,v 1.6 2005/01/24 14:47:04 morikawa Exp $

• Tag Name: $Name: $

• Change History:

1.1 Overview

Call hs94forcing mod .

1.2 Error Handling 1.3 Known Bugs 1.4 Note 1.5 Future Plans 1.6 Dependency

use type_mod,      only : STRING, DBKIND

1.7 Public Interface

private
public :: dycore_hs94forcing_init        ! subroutines
public :: dycore_hs94forcing             ! subroutines
public :: dycore_hs94forcing_end         ! subroutines

1.8 Procedure Interface

1.8.1 Initialize module and

hs94forcing mod を初期化する。

subroutine dycore_hs94forcing_init( Dims )

Dependency

use constants_mod  , only: constants_init
use spml_mod       , only: spml_init
use dycore_type_mod, only: DYCORE_DIMS
use dycore_grid_mod, only: dycore_grid_init, im, jm, km, nm
use hs94forcing_mod, only: hs94forcing_init
use dc_trace,        only: BeginSub, EndSub, DbgMessage

Input

type(DYCORE_DIMS), intent(in)   :: Dims   ! 次元データ全種

1.8.2 Calculate Heating and Dissipation for ”Held and Suarez(1994)”

subroutine dycore_hs94forcing( Vars_b, Vars_n, Vars_a )

Dependency

use dycore_type_mod, only: DYCORE_VARS, &
     &                     STRING, DBKIND, INTKIND
use dycore_time_mod, only: DelTime
use constants_mod  , only: R0
use spml_mod       , only: xya_wa, wa_Div_xya_xya, &
     &                     wa_LaplaInv_wa, wa_xya, &
     &                     xya_GradLat_wa, xya_GradLon_wa
use hs94forcing_mod, only: hs94forcing
use dc_trace       , only: BeginSub, EndSub, DbgMessage

Input

type(DYCORE_VARS), intent(in)   :: Vars_b ! 格子点データ全種 (t-
Δ t)
type(DYCORE_VARS), intent(in)   :: Vars_n ! 格子点データ全種 (t)

Output

type(DYCORE_VARS), intent(inout):: Vars_a ! 格子点データ全種 (t+
Δ t)

1.8.3 Terminate module

hs94forcing mod の終了処理を行なう。

subroutine dycore_hs94forcing_end

Dependency

use dycore_type_mod, only: STRING, DBKIND, INTKIND
use hs94forcing_mod, only: hs94forcing_end
use dc_trace,        only: BeginSub, EndSub, DbgMessage